Which of the following functions represents a cell phone bill if there is a monthly charge of 20 dollars plus 6 cents for each minute of use?
f(x) = 20 + 6x
f(x) = 20.06x
f(x) = 20 + 0.06x

Answer by stanbon(75887)

Which of the following functions represents a cell phone bill if there is a monthly charge of 20 dollars plus 6 cents for each minute of use?
---------
The 20 is dollars
Change 6 cents to 0.06 dollars
------
Let "x" be # of minutes of use.
----
Equation:
Cost = 20 + 0.06x dollars
Cost = 20.06x dollars
--------------------------
Cheers,
Stan H.
--------------------------
f(x) = 20 + 6x
f(x) = 20.06x
f(x) = 20 + 0.06x